# DistRL: An Asynchronous Distributed Reinforcement Learning Framework for On-Device Control Agents

### v2

@@ 6.2 Benchmarks and Baseline Methods
To comprehensively validate our approach, we utilize both the General and web shopping tasks for training and testing. Specifically, our training set is derived from enhanced online task instructions, which are composed of AitW Rawles et al. (2024b), AndroidWorld Rawles et al. (2024a), and expert-curated task sets. We fine-tune our model on this combined training set and evaluate performance on the corresponding test subsets derived from AitW. Our analysis focuses on training efficiency using the General Tasks, which include fundamental application operations and information retrieval tasks. Additionally, we assess the agent’s performance on both General Tasks and Web Shopping Tasks to evaluate its capability in handling domain-specific instructions, addressing the significant task distribution gap. Detailed descriptions of the datasets used are provided in Appendix A.5. Our baseline methodsA qualitative comparison among methods and more details can be found in Table 3 in Appendix A.5.1 include:
  • DigiRL Bai et al. (2024): The state-of-the-art framework prior to our work, which integrates RL fine-tuning with visual language models (VLMs) and provides a reproducible training process. We consider its both single and multi-machine settings in online mode.
  • AutoUI Zhan & Zhang (2023): A simple explorative mobile agent equipped with VLMs under supervised fine-tuning.
  • GPT-4V OpenAI et al. (2024) and Gemini 1.5 Pro Reid et al. (2024): Equipped with exploration drives-AppAgent Yang et al. (2023b) to facilitate learning from the environments.
@@ A.2.1 Evaluation with Auto-Evaluator
To generalize the evaluator across a wide range of tasks without manual rule definitions, we leverage a pre-trained LLM with appropriate prompting. The prompt is designed to instruct the LLM to act as an expert evaluator, i.e., Gemini-1.5-Pro Reid et al. (2024), in our practice. An example of such a prompt is provided below:
@@ A.4.1 Limitations of Traditional Methods
On-policy algorithms such as Proximal Policy Optimization (PPO Schulman et al. (2017)) and Advantage Actor-Critic (A2C Mnih (2016)) require synchronous data collection and updates, leading to inefficiencies in distributed and large-scale environments due to low sample efficiency and synchronization delays.
Standard off-policy methods like V-trace have been popular in distributed RL frameworks (e.g., IMPALA Espeholt et al. (2018)) but can be suboptimal when the divergence between the behavior policy $\mu$ and the target policy $\pi$ is either too small or too large due to clipping mechanisms.
@@ A.4.3 Hyperparameter Tuning for Distributed Prioritized Experience Replay
These ranges were selected based on insights from prior research Schaul (2015), Horgan et al. (2018) and preliminary experiments that indicated effective performance within these intervals.
@@ A.5.1 Baseline Methods
We evaluate proprietary vision-language models (VLMs), GPT-4V OpenAI et al. (2024) and Gemini 1.5 Pro Reid et al. (2024), using the AppAgent framework. By applying the prompt from  Yang et al. (2023b), we enable these models to interact effectively with the environment. We assessed the AppAgent Yang et al. (2023b) in a augmented prompting setting, where the agent explores the environment and gathers experience ahead of inference phase. This collected experience is appended to the test-time prompt, enhancing the model’s decision-making capabilities. Unlike learning-based approaches, these methods rely on advanced prompting strategies to accomplish tasks without extensive training. Additionally, our framework DistRL integrates GPT-2 Radford et al. (2019) for policy learning and ROBERTA-base Liu (2019) for the critic model We adopt the same baseline agent (AutoUI-driven agent) and VLMs models as used in DigiRL to validate the advantages of our fine-tuning framework and algorithm
@@ A.5.2 Training and Test Data
The dataset used in this work is based on the Android in the Wild (AitWRawles et al. (2024b) and AndroidWorld Rawles et al. (2024a) task set, with enhancements for practical use in fine-tuning agents to control mobile devices and interact with real-world applications. We trained two separate models using two distinct subsets for General Tasks and Web Shopping Tasks. Each model was trained on its corresponding training subset and evaluated on the respective test subset drawn from AitW.

### v3

@@ 6.2 Benchmarks and Baseline Methods
To comprehensively validate our approach, we utilize both the General and web shopping tasks for training and testing. Specifically, our training set is derived from enhanced online task instructions, which are composed of AitW (Rawles et al., 2024b), AndroidWorld (Rawles et al., 2024a), and expert-curated task sets. We fine-tune our model on this combined training set and evaluate performance on the corresponding test subsets derived from AitW. Our analysis focuses on training efficiency using the General Tasks, which include fundamental application operations and information retrieval tasks. Additionally, we assess the agent’s performance on both General Tasks and Web Shopping Tasks to evaluate its capability in handling domain-specific instructions, addressing the significant task distribution gap. Detailed descriptions of the datasets used are provided in Appendix A.5. Our baseline methodsA qualitative comparison among methods and more details can be found in Table 3 in Appendix A.5.1 include:
  • DigiRL (Bai et al., 2024): The state-of-the-art framework prior to our work, which integrates RL fine-tuning with visual language models (VLMs) and provides a reproducible training process. We consider its both single and multi-machine settings in online mode.
  • AutoUI (Zhan & Zhang, 2023): A simple explorative mobile agent equipped with VLMs under supervised fine-tuning.
  • GPT-4V (OpenAI et al., 2024) and Gemini 1.5 Pro (Reid et al., 2024): Equipped with exploration drives-AppAgent (Yang et al., 2023b) to facilitate learning from the environments.
@@ A.2.1 Evaluation with Auto-Evaluator
To generalize the evaluator across a wide range of tasks without manual rule definitions, we leverage a pre-trained LLM with appropriate prompting. The prompt is designed to instruct the LLM to act as an expert evaluator, i.e., Gemini-1.5-Pro (Reid et al., 2024), in our practice. An example of such a prompt is provided below:
@@ A.4.1 Limitations of Traditional Methods
Unlike model-free approaches used in applications such as Sun et al. (a; b), Sun & Wang (2022), Wang et al. (2018), Wang & Shi (2021), Wang & Yoneki (2024), Wang et al. (2024), Li et al. (2024b), Wang & Yoneki , Hadian et al. (2020), our RL algorithm for on-mobile training tasks requires customization and adaptation to meet specific task demands.On-policy algorithms such as Proximal Policy Optimization (PPO (Schulman et al., 2017)) and Advantage Actor-Critic (A2C (Mnih, 2016)) require synchronous data collection and updates, leading to inefficiencies in distributed and large-scale environments due to low sample efficiency and synchronization delays.
Standard off-policy methods like V-trace have been popular in distributed RL frameworks (e.g., IMPALA (Espeholt et al., 2018)) but can be suboptimal when the divergence between the behavior policy $\mu$ and the target policy $\pi$ is either too small or too large due to clipping mechanisms.
@@ A.4.3 Hyperparameter Tuning for Distributed Prioritized Experience Replay
These ranges were selected based on insights from prior research (Schaul, 2015, Horgan et al., 2018) and preliminary experiments that indicated effective performance within these intervals.
@@ A.5.1 Baseline Methods
We evaluate proprietary vision-language models (VLMs), GPT-4V (OpenAI et al., 2024) and Gemini 1.5 Pro (Reid et al., 2024), using the AppAgent framework. By applying the prompt from  (Yang et al., 2023b), we enable these models to interact effectively with the environment.
We assessed the AppAgent (Yang et al., 2023b) in a augmented prompting setting, where the agent explores the environment and gathers experience ahead of inference phase. This collected experience is appended to the test-time prompt, enhancing the model’s decision-making capabilities. Unlike learning-based approaches, these methods rely on advanced prompting strategies to accomplish tasks without extensive training. Additionally, our framework DistRL integrates GPT-2 (Radford et al., 2019) for policy learning and ROBERTA-base (Liu, 2019) for the critic model We adopt the same baseline agent (AutoUI-driven agent) and VLMs models as used in DigiRL to validate the advantages of our fine-tuning framework and algorithm
@@ A.5.2 Training and Test Data
The dataset used in this work is based on the Android in the Wild (AitW(Rawles et al., 2024b) and AndroidWorld (Rawles et al., 2024a) task set, with enhancements for practical use in fine-tuning agents to control mobile devices and interact with real-world applications. We trained two separate models using two distinct subsets for General Tasks and Web Shopping Tasks. Each model was trained on its corresponding training subset and evaluated on the respective test subset drawn from AitW.